Today, Maria's age is six times than Ana's age ten years ago. Five years from now, Ana's age will be the same as Maria's age ten years ago. How old is Maria
Answer by ankor@dixie-net.com(22740)

write an equation for each statement
Today, Maria's age is six times than Ana's age ten years ago.
m = 6(a-10)
distribute
m = 6a - 60
Five years from now, Ana's age will be the same as Maria's age ten years ago.
a + 5 = m - 10
a = m - 10 - 5
a = m - 15
or
m = a + 15
therefore
6a - 60 = a + 15
6a - a = 15 + 60
5a = 75
a = 75/5
a = 15 yrs is Ana's age
How old is Maria
m = 15 + 15
m = 30 yrs is Maria
:
:
Check this using the statement
"Five years from now, Ana's age will be the same as Maria's age ten years ago."
15 + 5 = 30 - 10